OpenStudy (anonymous):

Find the exact value of the expression:

OpenStudy (anonymous):

\[\sin^{-1} (-\sqrt2/2) \] The book says 3pi/4, why is the answer but 5pi/4?

OpenStudy (anonymous):

not*

OpenStudy (anonymous):

the book is definitely wrong because \(\large sin(\frac{3\pi}{4})=\frac{\sqrt2}{2} \) ...

OpenStudy (anonymous):

Sorry the book says -pi/4, but why would it not be something from a different quadrant where the y value is the same?

OpenStudy (anonymous):

oh... okay... the range of the inverse sine function is \( (-\frac{\pi}{2}, \frac{\pi}{2}) \)

OpenStudy (anonymous):

oh, so the other ones that correspond get ruled out then?

OpenStudy (anonymous):

yes....

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

yw...

OpenStudy (anonymous):

wait.. sorry.. the range is \(\ [-\frac{\pi}{2}, \frac{\pi}{2}] \) and NOT \(\ (-\frac{\pi}{2}, \frac{\pi}{2}) \) ...

OpenStudy (anonymous):

got it